### v1

@@ 5.1 Experimental Setup
Target Models and Datasets
We evaluate two representative open-source DLMs: Dream-7B-Instruct ye2025dream7b and LLaDA-8B-Instruct nie2025large. Our prompts are derived from Bespoke-Stratos-17k bespoke_stratos. Specifically, we use two publicly available datasets from the Hugging Face Hub lansechen_easy_2025; lansechen_hard_2025, retaining items whose prompt length is $\leq 512$ tokens. For LLaDA, we further augment its corpus with 7.57.5k additional math-style prompts sourced from the DParallel dataset chen2025dparallel. For each prompt, we generate teacher trajectories at temperatures $\{0.0,\,0.5\}$ (greedy / non-greedy), yielding 1515k generations for Dream and 3030k for LLaDA. Generation uses the bidirectional, block-wise teacher with generation length $L_{g}{=}256$, block size B=32B{=}32, and total steps N=256N{=}256. Additional details are provided in Appendix A.1.
@@
[…]
@@
Following established conventions wang2025d2f; wu2025fastdllm, we evaluate mathematical reasoning and code generation on GSM8K, GSM8K-CoT cobbe2021gsm8k, MATH hendrycks2021math, HumanEval chen2021humaneval, and MBPP austin2021mbpp. We use the lm_eval framework eval-harness and the Instruct task variants for the coding benchmarks. We report exact match scores on GSM8K, math-verify scores on MATH, and pass@1 results on HumanEval, HumanEval-Instruct, and MBPP-Instruct after standard post-processing. Coding tasks are evaluated zero-shot, and few-shot settings for math follow the original Dream and LLaDA papers. All efficiency measurements are taken on $4\times$ NVIDIA A100 (80 GB) GPUs with batch size $1$ under data parallelism. Latency, total steps, and generation length are reported as per-sample averages over the evaluation set.
@@
As references, we evaluate the original DLMs under their official inference settings. For LLaDA, we enable block-wise decoding. Dream does not support block-wise decoding, so we extend the baseline accordingly. Because our work targets two orthogonal axes of inference acceleration of (1) KV caching and (2) step reduction, we include three inference-time baselines that isolate or combine these factors: (i) dLLM-Cache liu2025dllmcache, which focuses on (1) via adaptive feature caching; (ii) Fast-dLLM (Parallel) wu2025fastdllm, which targets (2) via confidence thresholding; and (iii) Fast-dLLM (Parallel + Dual Cache) wu2025fastdllm, which combines (1) + (2) via approximate dual-cache KV caching. We omit D2F wang2025d2f here because it is trained for $L_{g}{=}512$ and reports on Dream-7B-Base rather than Dream-7B-Instruct. For autoregressive comparisons, we evaluate Qwen2.5-7B-Instruct qwen2024techreport alongside Dream and Llama-3.1-8B-Instruct dubey2024llama alongside LLaDA. For fair comparison, we use a fixed generation length of $L_{g}{=}256$, greedy decoding (temperature $0.0$), a block size of B=32B{=}32 for block-wise decoding, and a token-confidence threshold of $\tau_{\text{conf}}{=}0.9$. Further details are in Appendix A.3.
@@ 5.2 Main Results
We compare CDLM against vanilla DLMs, accelerated DLM variants, and autoregressive baselines, and summarize the quantitative results.
@@ 5.2.1 Results on CDLM–Dream
We attribute the drop on MATH to two factors. First, the small training mixture ($\sim$7.5k prompts) limits exposure to the advanced problem-solving skills required by MATH. Since the mixture is drawn from filtered subsets of Bespoke-Stratos-17k that largely consist of problems Qwen2.5 can already solve, the effective difficulty range is likely narrower than that of MATH. Second, CDLM is a fine-tuning-based method trained with teacher trajectories capped at 256 generation tokens. While sufficient for other benchmarks, this shorter thinking budget may be inadequate for complex multi-step reasoning. A natural next step would be to train with a 512-token budget, as D2F wang2025d2f does.
Latency.
@@
Throughput (TPS).
CDLM–Dream achieves the highest TPS on three out of four tasks, outperforming all baselines. On MBPP-Instruct, it jumps from 2.3 to 48.1 tokens/s (a $20.9\times$ increase over the naive model). The one exception is HumanEval-Instruct, where our TPS is lower than Fast-dLLM baselines because our effective generation length is much shorter (about 97 tokens vs. $\sim$200 for the baselines). This can depress our measured TPS and inflate baseline TPS. Crucially, despite shorter generations, our pass@1 is higher, indicating that CDLM–Dream emits fewer redundant tokens while maintaining or improving solution quality.
@@ 5.2.2 Results on CDLM–LLaDA
Figure 3: Throughput vs. AR (Dream). Tokens-per-second on GSM8K-CoT and MBPP-Instruct for Dream-7B-Instruct (naive), Qwen2.5-7B-Instruct (AR), and CDLM–Dream.
Figure 4: Throughput vs. AR (LLaDA). Tokens-per-second on GSM8K and HumanEval for LLaDA-8B-Instruct (naive), Llama3.1-8B-Instruct (AR), and CDLM–LLaDA.
@@ 5.2.3 Comparison with Autoregressive Models
In terms of throughput, CDLM improves naive DLMs by $3\times$$21\times$ and also surpasses equal-size autoregressive (AR) baselines. For Dream, we compare against Qwen2.5-7B-Instruct (Figure 3), where CDLM–Dream achieves $1.21\times$ higher throughput on GSM8K-CoT and $1.12\times$ on MBPP-Instruct. For LLaDA, we compare against Llama-3.1-8B-Instruct (Figure 4), where CDLM–LLaDA reaches $1.27\times$ the throughput on GSM8K and $4.17\times$ on HumanEval. Throughput here reflects both per-step cost and the number of refinement steps. A CDLM refinement step is more expensive than an AR step due to matrix-matrix multiplications for decoding. However, CDLM can finalize multiple tokens per iteration, reducing total steps and increasing effective throughput. For example, from Table 1, CDLM–Dream produces on average $107.4/44.1\approx 2.4$ tokens per step on GSM8K-CoT and $74.3/33.2\approx 2.2$ tokens per step on MBPP-Instruct.
@@ 5.3 Ablation Studies
Table 3: Ablation of loss weights. We vary $(w_{\text{distill}},w_{\text{cons}},w_{\text{dlm}})$ and report score (top, black) and steps to convergence (bottom, blue) on GSM8K (4-shot) and HumanEval-Instruct (0-shot) for CDLM–Dream.
@@ 5.3.1 Loss-Weight Composition
Removing the ground-truth masked-denoising term (rows 4 and 6) raises HumanEval but lowers GSM8K, indicating a loss of math reasoning ability. A small auxiliary weight helps preserve this ability, and we find that general coding performance can be recovered with longer training.
Based on these trends, we fix $w_{\text{distill}}{=}1.0$ and empirically choose $w_{\text{cons}}{=}0.5$ as a good balance between speed and quality. For the auxiliary DLM loss, we use $w_{\text{dlm}}{=}0.01$ for Dream and $w_{\text{dlm}}{=}0.1$ for LLaDA, since the DLM loss on LLaDA has a smaller absolute scale.
Table 4: Ablation of refinement steps. We report GSM8K results after naively truncating baseline DLM step counts to approximately match the step budgets used by CDLM–Dream and CDLM–LLaDA.
@@
5.3.2 Effective Step Reduction
We evaluate a naive step-truncation baseline by forcing the teacher DLMs to use a similar number of refinement steps as our CDLMs (Table 4). Because decoding is block-wise, the step count must be a multiple of the block factor ($L_{g}/B{=}8$). For Dream, we set the step budget to $48$ and measure GSM8K-CoT (8-shot) as in Table 1; for LLaDA, we set it to $56$ and measure GSM8K (4-shot) as in Table 2. Naively truncating the step budget, in other words, forcing a non-retrained DLM to finalize multiple tokens per iteration, leads to marked accuracy degradation. By contrast, CDLM maintains quality while operating with substantially fewer steps, underscoring that consistency training is necessary for stable multi-token refinement. Furthermore, with KV caching, our approach reduces latency by roughly half relative to the naive DLM at comparable iteration counts.
Table 5: Ablation of the confidence threshold. We sweep the token-level threshold $\tau_{\text{conf}}$ for CDLM–Dream on GSM8K-CoT (8-shot) and HumanEval-Instruct (0-shot).
GSM8K
-CoT
(8-shot)
0.95 42.7 2.5 78.8
0.90 51.7 2.1 78.8
0.85 57.7 1.9 78.4
HumanEval
-Instruct
(0-shot)
0.95 34.4 2.8 51.2
0.90 43.3 2.2 50.0
0.85 47.3 2.0 48.2
5.3.3 Token-Level Confidence Threshold
@@ 6 Discussion
Here, we discuss some complementary and future avenues for our work, as well as its limitations.
Inference-only accelerations.
Our method complements and can be combined with various inference-only (non-trainable) acceleration techniques. At inference time, we use only block-wise KV caching and confidence-based multi-token finalization, but more sophisticated mechanisms can be layered on top. For example, one can incorporate D2F’s inter-block parallelism to further reduce wall-clock latency wang2025d2f.
DLM as a drafter.
Speculative decoding employs a smaller draft model to propose token bursts that a larger verifier accepts or rejects in parallel leviathan2023fast. A naive diffusion language model (DLM) drafter yields limited benefit because standard DLMs require many refinement steps to produce coherent text. In contrast, a consistency-trained DLM (CDLM) can generate reasonable drafts in far fewer steps. Pairing a CDLM drafter with an autoregressive (AR) verifier (cf. christopher2024specdiff) is promising: the CDLM proposes multi-token candidates quickly, and the AR model supplies strong final validation.
@@
Training currently relies on offline, static trajectories. Despite careful dataset selection, the student may overfit to the teacher’s prior because supervision does not adapt during learning. For LLaDA, we observe that augmenting the training set with additional math-style prompts improves math performance by 2–3 percentage points while preserving other abilities, suggesting that scale and domain coverage are critical factors. Scaling beyond our current 15k-prompt corpus and broadening it to more diverse domains are straightforward directions for future work. Furthermore, CDLM’s performance is bounded by the teacher: a bidirectional DLM distilled into a block-causal student cannot exceed the teacher’s knowledge. Distilling from stronger DLM or AR teachers is a natural way to lift this ceiling. Additional discussion appears in Appendix B.
@@ 7 Conclusion
We presented CDLM, a training-based acceleration scheme that brings consistency modeling to DLMs. By enforcing within-block temporal consistency and fine-tuning a block-wise causal student, CDLM reduces refinement steps and enables natural KV caching. Across math and coding tasks, it yields faster inference, fewer steps, lower latency, and higher throughput, while maintaining competitive accuracy. Promising directions include scaling the distillation corpus, broadening domain coverage, and distilling from stronger teachers.
@@ Acknowledgements
[…]
Figure 6: Teacher trajectory and hidden-state buffer. Left: Block-wise decoding trajectory. Right: As each token is finalized, the teacher’s last hidden state at that position is written to a fixed-length buffer of size $L_{g}$; arrows indicate write indices across steps (toy example with B=3B{=}3, $L_{g}{=}6$).
@@ A.1 Dataset and Preprocessing
We use two datasets lansechen_easy_2025; lansechen_hard_2025, which are post-processed versions of Bespoke-Stratos 17k bespoke_stratos. We choose them because (i) D2F wang2025d2f, a representative fine-tuning-based DLM acceleration method, uses the same source, and (ii) they are filtered to a maximum length of 600 tokens and predominantly contain math word-style reasoning problems. For both Dream and LLaDA, we treat Qwen2.5-7B’s responses as ground-truth answers.
Dream is trained only on this Bespoke-derived subset. To strengthen math performance for LLaDA, we further augment its corpus with an additional 7.57.5k math-style prompts sampled from the DParallel dataset chen2025dparallel, matching the number of Bespoke-derived prompts. The released DParallel corpus does not include Qwen2.5-7B outputs, so we generate reference answers with Qwen2.5-7B for these prompts, yielding an LLaDA training set whose size is approximately $2\times$ that of Dream. After trying multiple datasets, we empirically observe that dataset selection is crucial; overly rigid formats (e.g., multiple-choice math) tend to hurt the model’s ability to generalize.
@@
In DLMs, temperature influences not only token choices but also the order in which tokens are revealed gong2025diffucoder. We therefore generate multiple trajectories per prompt at different temperatures. However, as shown in Figure 5, $\tau{=}1.0$ often destabilizes the reasoning chain and can yield incorrect conclusions. Consequently, we restrict trajectory collection to $\tau\in\{0.0,\,0.5\}$.

### v2

@@ 5.1 Experimental Setup
Target Models and Training Datasets
We consider two representative open-source DLMs: Dream-7B-Instruct Ye et al. (2025) and LLaDA-8B-Instruct Nie et al. (2025b). We derive 7.5k prompts from Bespoke-Stratos-17k Labs (2025), filtering to those with prompt length $\leq 512$ tokens. For LLaDA, we further augment the corpus with 7.5k math-style prompts from DParallel Chen et al. (2025), using the same length budget. When generating teacher trajectories, we fix the generation length to $L_{g}{=}256$ and the block size to B=32B{=}32. Additional details are provided in Appendix A.1.
@@
We fine-tune all models with LoRA Hu et al. (2022) applied to both attention and MLP modules, using a block-wise causal mask (Figure 2) with B=32B{=}32 and $L_{g}{=}256$. End-to-end training takes approximately 8 hours for CDLM–Dream and 16 hours for CDLM–LLaDA on $4\times$ NVIDIA A100 (80GB) GPUs. Additional details are in Appendix A.2.
@@
Following established conventions Wang et al. (2025); Wu et al. (2025b), we evaluate mathematical reasoning and code generation on GSM8K, GSM8K-CoT Cobbe et al. (2021), MATH Hendrycks et al. (2021), HumanEval Chen et al. (2021a), and MBPP Austin et al. (2021b). Coding tasks are evaluated zero-shot, and few-shot settings for math follow the original Dream and LLaDA papers. All efficiency measurements are taken on $4\times$ NVIDIA A100 (80 GB) GPUs with batch size $1$ under data parallelism. Latency, total steps, and generation length are reported as per-sample averages over the evaluation set. Additional evaluation details are provided in Appendix A.3.
@@
As references, we evaluate the original DLMs using block-wise decoding under their official inference settings. Because our work targets two orthogonal axes of inference acceleration, (1) KV caching and (2) step reduction, we include three inference-time baselines: (i) dLLM-Cache Liu et al. (2025b), which focuses on (1) via adaptive feature caching; (ii) Fast-dLLM (Parallel) Wu et al. (2025b), which targets (2) via confidence thresholding; and (iii) Fast-dLLM (Parallel + Dual Cache) Wu et al. (2025b), which combines (1) + (2) via approximate dual-cache KV caching. We omit D2F Wang et al. (2025) here because it is trained for $L_{g}{=}512$ and reports on Dream-7B-Base rather than Dream-7B-Instruct. For autoregressive comparisons, we evaluate Qwen2.5-7B-Instruct Qwen et al. (2024) alongside Dream and Llama-3.1-8B-Instruct Grattafiori et al. (2024) alongside LLaDA. For fair comparison across methods, we fix $L_{g}{=}256$, use greedy decoding (temperature $0.0$), apply the same B=32B{=}32, and set the token-confidence threshold to $\tau_{\text{conf}}{=}0.9$. Further details are in Appendix A.3.
@@ 5.2 Main Results
We compare CDLM against vanilla DLMs, accelerated DLM variants, and autoregressive baselines.
@@ 5.2.1 Results on CDLM–Dream
We attribute the accuracy drop on MATH to two factors. First, the small training mixture ($\sim$7.5k prompts) provides limited exposure to the advanced reasoning patterns required by MATH. The data are drawn from filtered subsets that largely contain problems already solvable by Qwen2.5, resulting in a narrower effective difficulty range. Second, teacher trajectories are capped at 256 generation tokens during training. This shorter reasoning budget may be inadequate for complex multi-step problems. A natural next step is to train with a longer budget (e.g., 512 tokens), as in D2F Wang et al. (2025).
Latency and Throughput (TPS).
@@
CDLM–Dream attains the highest TPS on three out of four benchmarks. The sole exception is HumanEval-Instruct, where TPS is lower than Fast-dLLM baselines due to a substantially shorter effective generation length (97 vs. $\sim$200). Differences in decoding dynamics can arise because CDLM is strictly block-causal and cannot attend to future blocks, lacking explicit budget awareness, whereas other baselines retain full bidirectional attention that preserves a form of global budget awareness. Despite shorter generations, CDLM–Dream achieves higher pass@1, indicating that it can emit fewer tokens while preserving solution quality.
@@ 5.2.2 Results on CDLM–LLaDA
Figure 3: Throughput comparison across benchmarks. Tokens per second on GSM8K, MBPP, and HumanEval for Dream-7B-Instruct and LLaDA-8B-Instruct under naive diffusion decoding, autoregressive (AR) baselines, and CDLM.
@@ 5.2.3 Comparison with Autoregressive Models
CDLM improves naive DLMs in throughput by $3\times$$21\times$ and also surpasses equal-size autoregressive (AR) baselines (Figure 3). CDLM–Dream achieves $1.2\times$ and $1.1\times$ higher throughput than its AR baseline on GSM8K-CoT (8-shot) and MBPP-Instruct (0-shot), respectively, while CDLM–LLaDA attains $1.3\times$ and $4.2\times$ higher throughput on GSM8K (4-shot) and HumanEval (0-shot). Throughput here reflects both per-step cost and the number of refinement steps. A CDLM refinement step is more expensive than an AR step due to matrix-matrix multiplications for decoding. However, CDLM can finalize multiple tokens per iteration, reducing total steps and increasing effective throughput. For example, from Table 1, CDLM–Dream produces on average $107.4/44.1\approx 2.4$ tokens per step on GSM8K-CoT and $74.3/33.2\approx 2.2$ tokens per step on MBPP-Instruct.
@@ 5.3 Ablation Studies
Table 3: Ablation of loss weights. Effects of varying $(w_{\text{distill}},w_{\text{cons}},w_{\text{dlm}})$ on score (black) and steps to convergence (blue) on GSM8K and HumanEval-Instruct for CDLM–Dream.
@@ 5.3.1 Loss-Weight Composition
Removing the ground-truth masked-denoising term (rows 4 and 6) raises HumanEval but lowers GSM8K, indicating a loss of math reasoning ability. A small auxiliary weight helps preserve this ability, and we find that general coding performance can be recovered with longer training. Accordingly, we use $w_{\text{dlm}}{=}0.01$ for Dream and $w_{\text{dlm}}{=}0.1$ for LLaDA, as the DLM loss on LLaDA has a smaller absolute scale.
5.3.2 Effective Step Reduction
We evaluate a naive step-truncation baseline by forcing the teacher DLMs to use a similar number of refinement steps as our CDLMs (Table 4). Because decoding is block-wise, the step count must be a multiple of the block factor ($L_{g}/B{=}8$). For Dream, we set the step budget to $48$ and measure GSM8K-CoT (8-shot) as in Table 1; for LLaDA, we set it to $56$ and measure GSM8K (4-shot) as in Table 2. Naively truncating the step budget, in other words, forcing a non-retrained DLM to finalize multiple tokens per iteration, leads to marked accuracy degradation. By contrast, CDLM maintains quality at comparable step counts, underscoring that consistency training is necessary for stable multi-token refinement. Furthermore, with KV caching, our approach reduces latency by roughly half relative to the naive DLM at comparable iteration counts.
Table 4: Ablation of refinement steps. GSM8K results obtained by naively truncating the baseline DLM step counts to match the step budgets used by CDLM–Dream and CDLM–LLaDA.
@@ 5.3.2 Effective Step Reduction
Figure 4: Arithmetic intensity across batch sizes. Arithmetic Intensity (AI) during decoding as a function of batch size ($\mathrm{bs}\in\{1,2,4,8,16,32,128\}$) for autoregressive (AR) models, vanilla DLMs, and block-wise DLMs (CDLM). The dashed line indicates the theoretical ridge point, separating memory-bound (red) and compute-bound (blue) regimes.
5.4 System-Level Scalability via Arithmetic Intensity
To understand CDLM’s hardware-utilization characteristics, we analytically model the inference for autoregressive (AR) models, vanilla DLMs, and block-wise DLMs with block sizes $B\in\{4,16,32\}$. Building on prior modeling frameworks Tiwari et al. (2025); Kim et al. (2025), we parameterize the AR baseline following the LLaMA-3.1 Grattafiori et al. (2024) configuration, while vanilla and block-wise DLMs follow the LLaDA Nie et al. (2025b) configuration. We examine arithmetic intensity (AI), the ratio of computation to memory traffic (FLOPs per byte moved), to diagnose whether inference is memory- or compute-bound and to characterize batch-level scalability. Concretely, we consider a prompt length $L_{p}{=}512$ and generation length $L_{g}{=}256$ to match Section 5.2, and track AI as the batch size ($\mathrm{bs}$) increases from 1 to 128 (Figure 4). To focus on decoding behavior, we exclude one-time prefill costs (KV-cache initialization) and ignore memory-capacity constraints (e.g., out-of-memory).
Autoregressive (AR) Models.
[…]
Vanilla DLMs.
[…]
Block-wise DLMs (CDLM).
Block-wise DLMs operate in an intermediate regime. At $\mathrm{bs}=1$, their AI (4.0, 15.8, 31.1 for $B\in\{4,16,32\}$) exceeds AR decoding but remains below vanilla DLMs. This gain stems from intra-block parallelism: each step processes $B$ tokens under comparable memory traffic as a single-token AR step, effectively amortizing weight loads across the block and scaling AI by nearly $B$. The AI crosses the ridge point at relatively small batch sizes (e.g., B=32B{=}32 at $\mathrm{bs}\approx 8$ and B=16B{=}16 at $\mathrm{bs}\approx 16$), allowing block-wise DLMs to better utilize available compute in small-batch settings. At higher batch sizes, however, the marginal AI gain diminishes as $\mathrm{bs}$ increases, since block-wise DLMs already exploit both batch- and block-level parallelism.
Overall, this analysis shows that CDLM-like block-wise DLMs sit between the memory-bound regime of AR models and the compute-saturated regime of full-attention DLMs. This balanced operating point explains CDLM’s superior throughput and scalability in small-batch inference. A detailed roofline analysis with AI and corresponding simulated performance is provided in Appendix B.4.
@@ 6 Discussion
Here, we discuss key trade-offs, potential extensions, and limitations of our approach.
Expressiveness vs. efficiency trade-off
Full bidirectional attention in DLMs requires recomputing $\mathcal{O}(L^{2})$ attention at every denoising step, making inference highly compute-intensive and impractical at scale. CDLM enables exact KV caching to mitigate this cost while preserving bidirectional context within each block, retaining local refinement capabilities such as infilling inside the current block. We further believe that the mild left-to-right inductive bias introduced by block-wise decoding aligns well with reasoning-style generation.
Inference-only acceleration techniques
Our method is orthogonal to and can be combined with various inference-only (training-free) techniques. At inference time, we only use block-wise KV caching and confidence-based multi-token finalization, but more sophisticated mechanisms can be layered on top. For example, incorporating D2F Wang et al. (2025)’s inter-block parallelism could further reduce wall-clock latency.
Scaling with stronger DLM backbones
CDLM is a post-training recipe that can be applied to any block-diffusion models to accelerate convergence, and its effectiveness is expected to grow as stronger DLMs become available. Recent models such as SDAR Cheng et al. (2025) are reported to outperform Dream and LLaDA and have been released at multiple scales. A promising direction is to generate trajectories from large-scale DLM teachers (e.g., 30B) and train mid-scale students (e.g., 8B) using the CDLM post-training recipe.
@@
Training currently relies on offline, static trajectories that are predominantly math-focused. Despite careful dataset selection, the student may overfit to the teacher’s prior and domain, potentially limiting generalization beyond the training distribution. For LLaDA, we observe that augmenting the training set with additional math-style prompts improves math performance by 2–3 percentage points while preserving coding ability, suggesting that scaling and diversifying the corpus are straightforward directions for future work. Furthermore, CDLM’s performance is ultimately bounded by the teacher: a bidirectional DLM distilled into a block-causal student cannot exceed the teacher’s knowledge. Distilling from stronger DLMs or autoregressive teachers is a natural way to lift this ceiling. Additional discussion appears in Appendix C.
@@ 7 Conclusion
We presented CDLM, a training-based acceleration scheme that brings consistency modeling to DLMs. By enforcing within-block temporal consistency and fine-tuning a block-wise causal student, CDLM reduces refinement steps and enables exact KV caching. Across math and coding tasks, it yields faster inference, fewer steps, lower latency, and higher throughput while maintaining competitive accuracy. Promising directions include scaling the distillation corpus, broadening domain coverage, and distilling from stronger teachers.
@@ A.1 Dataset and Preprocessing
We use two publicly available datasets from the Hugging Face Hub Lansechen (2025a; b), which are post-processed versions of Bespoke-Stratos 17k Labs (2025). We choose them because (i) D2F Wang et al. (2025), a representative fine-tuning-based DLM acceleration method, uses the same source, and (ii) they are filtered to a maximum length of 600 tokens and predominantly contain math word-style reasoning problems. For both Dream and LLaDA, we treat Qwen2.5-7B’s responses as ground-truth answers.
Dream is trained only on this Bespoke-derived subset. To strengthen math performance for LLaDA, we further augment its corpus with an additional 7.57.5k math-style prompts sampled from the DParallel dataset Chen et al. (2025). The released DParallel corpus does not include Qwen2.5-7B outputs, so we generate reference answers with Qwen2.5-7B for these prompts, yielding an LLaDA training set whose size is approximately $2\times$ that of Dream. After trying multiple datasets, we empirically observe that dataset selection is crucial; overly rigid formats (e.g., multiple-choice math) tend to hurt the model’s ability to generalize.
Preprocessing and Generation Configuration.
We retain prompts whose input length is at most $512$ tokens. For each prompt, we generate teacher trajectories using the bidirectional, block-wise DLM, with generation length $L_{g}{=}256$, block size B=32B{=}32, total diffusion steps N=256N{=}256, and prompts left-padded to a length of 512 tokens. Ground-truth responses are truncated or padded to a length of 256 tokens.
@@
Effective training benefits from trajectories that are both diverse and reliable. In DLMs, temperature influences not only token choices but also the order in which tokens are revealed Gong et al. (2025b), enabling additional diversity in the generated trajectories. We therefore generate multiple trajectories per prompt at different temperatures. As shown in Figure 5, using temperature $\tau{=}1.0$ can destabilize the reasoning chain and lead to incorrect conclusions. Consequently, we restrict trajectory collection to $\tau\in\{0.0,\,0.5\}$.